S-1+2+3+4+...+1000? S-1+3+9+27+81+...-? S-25/1.08+25/1.08 2+25/1.08A3+...+25/1.08 20-? S-1+4+8+16+32+64+...+2A100-? S-30/1.07 2+30/1.07A3+.. +30/1.07A40-? S-30/1.07 2+30/1.07A3+30/1.07A4+...-1? 142 1+k None of the answers are correct 3t and kg TdE None of the answers are correct e-g
Answer #1

(a) S = 1+2+3+4+....+1000

This is an arithmetic progression

S = (n/2)[2a + (n-1)d] = (1000/2)[2 + (1000-1)*1] = 500500

(b) S = 1+3+9+27+ ...

This is a Geometric progression

S = 30 + 31 + 32 + 33 + 34 + ...

S = a(rn-1)/(r-1) = 1*(3n-1)/(3-1) = (3n-1)/2

(c) S = 25/1.08 + 25/1.08^2 + ...

This is a Geometric Progression

S = a(rn-1)/(r-1) = 25*((1/1.08)n-1)/((1/1.08)-1) = 337.5(1.08n-1)

(d) S = 1+4+8+16+...2^100

This is a geometric Progression

S = (20 + 21 + 22 + 23 +... 2100) - 21 = (2100 - 1)/(2-1) - 21 = 2100 - 3